Discount Rental Cars charges a daily fee plus a mileage fee for renting its cars. Barney was charged $145.00 for 3 days and 310 miles, while Mary was charged $250.00 for 5 days and 600 miles. What does Discount Rental Cars charge per day and per mile?please show me the steps that you did to get to the answer.
Answer by ikleyn(52850)

.

Let "d" be the charge per day and "m" be the charge per mile.


Then from the condition, you have these two equations


    3d + 310m = 145.00  dollars    (1)   (how much Barney was charged)

    5d + 600m = 250.00  dollars    (2)   (how much Mary   was charged)


At this point, the setup is just completed.

This part of the solution is rather simple.

It is quite obvious and usually does not create any difficulties.


Thus you have this system of equations, and now your task is to solve it.


There are different methods of solution : the Substitution method;  the Elimination method and the Cramer's rule (using determinants).


I will use the Elimination method. 
For It, I multiply equation (1) by 5 (both sides) and multiply equation (2) by 3.


You will get modified equivalent system


    15d + 1550m = 725.00      (3)  

    15d + 1800m = 750.00      (4)   (how much Barney was charged)


Notice the the coefficients at "d" are equal in both equation.


Next step, subtract equation (3) from equation (4).  The terms "15d" in both equations will cancel each other, 
and you will get a single equation for one unknown "m" :


    1800m - 1550m = 750.00 - 725,    or

    250m          = 25,

       m          = 25/250 = 0.1.


Thus you just found the unknown "m" : it is 0.1  (dollar per mile).


To get the second unknown, simply substitute this value m= 0.1 into either of the two original equations.

I will substitute into equation (1). Then you get


    3d + 310*0.1 = 145,

    2d + 31      = 145,

    3d           = 145 - 31 = 114,

     d           = 114/3 = 38.


Answer.  The company charges  38 dollars per day and 0.1 dollars (= 10 cents) per mile.

The problem is just solved.

The last step is to check the answer.
For it, substitute the found values into the original equations.
I expect that you do it on your own.

The solution is completed.

------------------

If you are a novice in solving word problems on two equations in two unknown, look into my lessons in this site
    - Oranges and grapefruits
    - Using systems of equations to solve problems on tickets
    - Three methods for solving standard (typical) problems on tickets
    - Using systems of equations to solve problems on shares
    - Using systems of equations to solve problems on investment
    - Two mechanics work on a car
    - The Robinson family and the Sanders family each used their sprinklers last summer
    - Roses and vilolets
    - Counting calories and grams of fat in combined food
    - A theater group made appearances in two cities
    - HOW TO algebreze and solve this problem on 2 equations in 2 unknowns
    - One unusual problem to solve using system of two equations
    - Solving mentally word problems on two equations in two unknowns
    - OVERVIEW of lessons on solving systems of two linear equations in two unknowns

You will find there many typical word problems with detailed solutions,  clearly explained and demonstrating different solution methods.

Consider these lessons as your textbook,  problems book with solutions,  guide,  tutorials and  (free of charge)  home teacher.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Systems of two linear equations in two unknowns".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.